look at the partial differential equation
xux - tut= u

show that u( x , t) = xf( xt) is a general solution , where f is any differentiable function
1
Expert's answer
2012-12-13T08:12:35-0500
X, T - variable
()x, ()t, ()' - derivative
X*Ux - T*Ut = U
U (X, T) = X* F (X*T)
X* (X* F (X*T))x - T* (X* F (X*T))t = X* F (X*T)
X* (F (X*T) + X*F'(X*T)*T) - T* (X* F'(X*T)* X) = X* F (X*T)
X*F (X*T) +{X^2*F'(X*T)*T - T* X^2* F'(X*T)} = X* F (X*T)
X*F (X*T)=X*F (X*T)